WARM, DENSE GAS NEAR THE MASSIVE PROTOSTAR AFGL 2136 IRS 1 AS REVEALED BY ABSORP-TION FROM THE ν_{1}, ν_{2}, AND ν_{3} BANDS OF WATER

\underline{NICK INDRIOLO}, DAVID A. NEUFELD, Department of Physics & Astronomy, Johns Hopkins University, Baltimore, MD 21218; ANDREAS SEIFAHRT, Department of Astronomy and Astrophysics, University of Chicago, Chicago, IL 60637; MATT J. RICHTER, Department of Physics, University of California Davis, Davis, CA 95616.

We have identified absorption features due to 47 different ro-vibrational transitions of the ν_{1} and ν_{3} fundamental and ν_{2} overtone bands of H_{2}O between 2.468 μm and 2.561 μm toward the massive protostar AFGL 2136 IRS 1, and determined column densities in each rotational level. Analysis of the relative level populations indicates the absorption arises in warm (T ≈ 500 K), very dense (n_{H} > 10^{10} cm^{−3}) gas in local thermodynamic equilibrium, with a total water column density of N(H_{2}O) ≈ 10^{19} cm^{−2}. This gas must be very close to the central protostar, either in the inner envelope or a circumstellar disk/torus. Our findings are consistent with results inferred from 4.7 μm CO observations,^{a} 6 μm H_{2}O observations,^{b}, and our own 2.5 μm HF observations.^{c} This study represents the first extensive use of water vapor absorption lines in the near infrared, and we examine the possibility of using such observations to derive physical parameters in the gas surrounding other protostars.

aMitchell, G. F., et al. 1990, ApJ, 363, 554

bBoonman, A. M. S., & van Dishoeck, E. F. 2003, A&A, 403, 1003

cIndriolo, N., Neufeld, D. A., Seifahrt, A., & Richter, M. J. 2013, ApJ, 764, 188